Operations Report – February 4th

Crew 188 Operational Report 04/02/2018

Sol 7

Zac Trolley

Non-Nominal Systems: electrical system, electrical cords

Notes on non-nominal systems: See individual systems at the end of the report.

Generator (hours run): 16.9h

Solar: SOC 80%

Diesel: 75%

Propane: 61 psi

Ethanol Free Gasoline (5 Gallon containers for ATV): 8 Gallons

Water (trailer): 0 Gallons

Water (static): 170 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Meter 130739.8

Toilet tank emptied: No

ATVs Used: None

Oil Added: None

ATV Fuel Used: None

# Hours the ATVs were Used today: 0

Notes on ATVs: none

Deimos rover used: No

Hours: Hours: 108.1

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours: 18.0h

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Opportunity rover used: No

Hours: 13.5

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Curiosity rover used: No

Hours: 11.3h

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

HabCar: Not used

General notes and comments: We are rotating the Rovers with the one plug in we have, and that should be sufficient if we are taking 2 rovers at a time per EVA.

It would seem that the electrical generation isn’t being regulated that well. The diesel and solar power systems will vary between 120V and 140V. No known reason why this is happening.

During an EVA, a plastic handle from the passenger side roll cage on Demos came off. We have the handle, but no hardware to re-attach it.

Generator was turned off at 10:36

Generator was turned on at 17:28

I’ve found two extension cords that have been damaged and need to be replaced. If there are 3 prong plugs available, I can repair the cords. Please advise.

Summary of internet: Internet connectivity continues to be a problem even with the data purchased. Devices are having troubles connecting and this is limiting the crew’s ability to monitor the CapCom emails

Summary of suits and radios: The clasp on Suit 1 was bent back to place, it’s as good as new.

Suit 2 was repaired, it was a cold solder joint, as well as the connection not being crimped. I expertly re-attached the connector and the suit is operational

Summary of Hab operations: Nominal

Summary of EVA operations: In the warm weather the fogging issue isn’t as pronounced. We are hoping for warm weather to maintain visual accuracy in the helmets.

Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.

Summary of Science Dome operations: The science dome is awaiting a portable heater as per the MDRS Director.

Summary of RAM operations: Not Operational.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support: We could use another bag of toilet deodorizer, as we only have one pack left.

Copyright © The Mars Society. All rights reserved. | Main Site